OpenStudy (tia_sweetheart17):

solve for x: -2(x-90)+4=22

OpenStudy (anonymous):

Do you have any ideas?

OpenStudy (amonoconnor):

-2(x - 9) + 4 = 22 -4 ________________ -2(x - 9) = 18 _____________ -2 X - 9 = (-9) +9 __________ X = 0 :)

OpenStudy (amonoconnor):

Does this algebraic process make sense?